A garden is 8 feet wide by 16 feet long. Make a scale drawing of the garden that has a scale of 1/4 in. = 2ft.
Answer by checkley71(8403)

8X16=(8/2)X(16/2)=(4X8)*1/4=1x2 INCHES IS THE SCALED DOWN DRAWING FOR THE 8X16 GARDEN USING 1/4 IN=2FT.